a gas has a density of 4.32 g/L. What volume in liters would 6.17 g of this gas occupy?
1
Expert's answer
2017-10-02T11:41:07-0400
V=m/ρ
V=6.17/4.32=1.43 (l)

Answer: 1.43 l.